OpenStudy (anonymous):

Prove that lim as x approaches 2 (1/itionx) = 1/2...is there anotehr way besides direct substitution

OpenStudy (anonymous):

lim x approaches 2 (1/x) = 1/x

OpenStudy (anonymous):

i want to prove it w. epsiolon and delta

OpenStudy (anonymous):

we can probably do this if you like

OpenStudy (anonymous):

no problem

OpenStudy (anonymous):

first i assume you mean \[\lim_{x\rightarrow 2}\frac{1}{x}=\frac{1}{2}\]

OpenStudy (anonymous):

so you have to show that given any epsilon greater than zero there will be some delta (which you write in terms of epsilon) so that if \[|x-2|<\delta\] then \[|\frac{1}{x}-\frac{1}{2}|<\epsilon\]

OpenStudy (anonymous):

as usual we work backwards to see what we need.

OpenStudy (anonymous):

\[|\frac{1}{x}-\frac{1}{2}|<\epsilon\] \[|\frac{2-x}{2x}|<\epsilon\] \[\frac{|2-x|}{|2x|}<\epsilon\] and now we see that we are almost there because \[|x-2|=|2-x|\] and that is the one we have control over

OpenStudy (anonymous):

the only thing we don't have a bound for is \[|2x|\] so we do the usual trick and make a bound for \[|x-2|\] say make sure that \[|x-2|<1\] which we are allowed since we control that term

OpenStudy (anonymous):

if \[|x-2|<1\] then \[-1<x-2<1\] so \[1<x<3\] and now we know that the smallest the denominator can be is 1

OpenStudy (anonymous):

or rather 2, since it is \[|2x|\] so we are in good shape now

OpenStudy (anonymous):

b4 we move on how is |x−2|=|2−x|

OpenStudy (anonymous):

if the smallest the denominator can be is 2, then \[\frac{|2-x|}{|2x|}\leq2|x-2|\] so make sure that and so if we make \[|x-2|<\frac{\epsilon}{2}\] then \[|\frac{2-x}{2x}|\leq 2|x-2|<2\frac{\epsilon}{2}=\epsilon\]and we are done

OpenStudy (anonymous):

it is always the case that \[|b-a|=|a-b|\]

OpenStudy (anonymous):

you can try it with numbers to see, but this says the distance between a and b is the same as the distance between b and a

OpenStudy (anonymous):

oh ok..i see cuz its absoute value

OpenStudy (anonymous):

right

OpenStudy (anonymous):

so if it equals epsilon at te hend taht means teh limit is proved or taht it actually esists?

OpenStudy (anonymous):

i think i made a mistake actually but you can always err on the side of caution. in fact if |2x|>2 then \[\frac{|2-x|}{|2x|}<\frac{1}{2}|x-2|\] so we could have picked a different delta, but as long as we get something smaller that is ok

OpenStudy (anonymous):

yes so if you can arrange it so that |x-a|< something insures that |f(x)-L|< epsilon, then you are done